# DeFT: Decoding with Flash Tree-Attention for Efficient Tree-structured LLM Inference

### v3

@@ A.3 Discussion of Concurrent Works
There are some concurrent works (Athiwaratkun et al., 2024; Ye et al., 2024; Juravsky et al., 2024) in attention algorithm design for single-context large-batch sampling, where the goal is to generate multiple sequences from a single context(e.g. system prompt or few-shot examples), which is a special case of tree-based decoding with a depth of 1. The design of their algorithms are based on this feature, which means they can not suit well in attention calculation of a tree with more than two levels of prefixes with efficiency.
@@
Table 9: Comparison among DeFT and concurrent works in single-context large-batch sampling scenarios, including Chunk-Attention (Ye et al., 2024), Hygragen (Juravsky et al., 2024) and Bifurcated-Attention (Athiwaratkun et al., 2024). More $\star$ means more balanced workloads after tree split, which also shows how insensitive the acceleration is to the tree topology.
@@ A.5 Analysis: IO Complexity of DeFT
Figure 11: (Supplementary to Figure 3) QKV partitioning of Tree Attention (Cai et al., 2024; Miao et al., 2023) and memory access. Tree Attention-Medusa (Cai et al., 2024) partitions QKV by General Matrix Multiply(GEMM) in PyTorch. Tree Attention-SpecInfer (Miao et al., 2023) adopts Q-Guided Grouping.
@@ A.6 Discussion of Workloads Generation
  • We can utilize decoding trees from real tasks as a benchmark within a unified system, enabling fair comparison of different attention algorithms in terms of wall-clock time performance. This comparison is possible despite the algorithms being based on distinct systems, such as variations in memory management implementations for their kernels.
@@ A.7 Additional Results
The details of attention latency and IO comparison among DeFT and baselines are in Table 5 and Table 17, respectively.
@@
Table 16: Average attention latency (in seconds) for tree-based decoding and its impact on end-to-end latency. Here, $b$ represents the tree width, and $t$ denotes the token tree size (i.e., the number of tree-structured queries). Attention Speedup over the best attention refers to the speedup of DeFT-Flatten compared to the best baseline (typically Tree Attention-Medusa) in attention calculation. Speedup over the best wall-clock time refers to the speedup of DeFT-Flatten compared to the best best wall-clock-latency baseline (Radix Attention). Attention Speedup over the best wall-clock indicates the attention speedup of DeFT-Flatten over the best wall-clock-latency baseline (Radix Attention). $\star$ denotes out-of-memory (OOM) errors for the A100 80GB GPU. For more details on end-to-end latency, see Table 5.
@@
Tree Attention-Medusa 3.93 7.51 9.57 38.64 29.10 2.62 3.96 18.05 26.31 41.10 68.28
Paged Radix Attention 5.99 7.30 9.96 39.37 24.69 3.11 5.13 32.60 54.57 109.39 212.29
DeFT-Flatten . 3.47 4.07 5.87 28.41 21.45 2.57 3.83 12.68 18.18 29.97 55.58
Attention Speedup over the best attention. 1.13$\times$ 1.63$\times$ 1.70$\times$ 1.36$\times$ 1.15$\times$ 1.02$\times$ 1.03$\times$ 1.42$\times$ 1.45×1.45$\times$1.45 × 1.37×1.37$\times$1.37 × 1.22×1.$2$
Attention Speedup over the best wall-clock 1.73$\times$ 1.63$\times$ 1.70$\times$ 1.39$\times$ 1.15$\times$ 1.21$\times$ 1.34$\times$ 2.57$\times$ 3.00×3.$0$ 3.64×3.64$\times$3.64 × 3.82×3.82$\times$3.82 ×
Speedup over the best wall-clock 1.24$\times$ 1.28$\times$ 1.33$\times$ 1.10$\times$ 1.03$\times$ 1.03$\times$ 1.05$\times$ 1.43$\times$ 1.70×1.70$\times$1.70 × $2\times2$ 2.52×2.52$\times$2.52 ×
Table 17: Average end-to-end IO (TB). Data format is Left/Right: (Left) KV Cache IO; (Right) partial results IO, including ${\bm{Q}}{\bm{K}}^{T}$,${\bm{Q}}{\bm{K}}^{\top}/{s_{c}}$, Mask $M$, ${\bm{M}}+{\bm{Q}}{\bm{K}}^{\top}/{s_{c}}$ and Softmax. $b$ means tree width. $t$ denotes the token tree size (i.e., the number of tree-structured queries)$\star$ means out of memory for A100 80GB.
@@
We observe that the effectiveness of attention speedup varies with different decoding tree topologies. Considering the simplest tree structure, a prompt with several suffixes—given a prompt that is not very short, one of the most important factors for speedup is the extent to which we can reuse its KV cache IO. This can be measured by the width of the tree. More specifically, it is determined by the number of queries per iteration. Therefore, we fix the prompt length at 4000 and vary the width of the decoding tree in few-shot prompting (which also indicates how many requests share the same prompt). Then, as shown in Figure 14, we evaluate DeFT-Flatten with the best baseline in attention calculation– Tree Attention-Medusa (Cai et al., 2024) (Medusa-Attn in the figure), as well as the best baseline in wall-clock time– Radix Attention (Zheng et al., 2023), for the per-iteration latency over time. We have the following observations:
@@
    Since the tree topology consists of a fixed prefix with several suffixes, a larger tree width allows the prompt prefix’s KV cache to be reused more frequently during IO. This leads to a more significant end-to-end speedup—1.24×1.24$\times$1.24 × with a width of $w=20$, and 1.33×1.$3$ with a width of $w=50$—compared to Radix Attention.
@@
Figure 14: Per iteration latency for few-shot prompting tasks with different tree width. e2e means end-to-end result, while Attn means only the attention overhead.
@@
In Table 6 and Table 8 of Section 4.4, we show the influence of prompt length and model size, individually. Here, we present the ablation study of both model size and prompt length for sorting task, as shown in Table 18. With a longer prompt, DeFT shows more pronounced speedup in the same model, since the attention overhead is proportional to the token count in the decoding tree, while the FFN overhead remains nearly constant for the same model. For a fixed prompt length, with the larger model—Codellama-34B, DeFT-Flatten achieves slightly reduced but still significant (up to 1.28x) end-to-end speedup. The performance reduction is attributed to a lower A/F-LR, as the FFN overhead is greater in larger models.
Table 18: [Ablation Study of Model Size and Prompt Length] Comparison of wall-clock time speedup and Attention/FFN latency ratio (A/F-LR) between DeFT and Radix Attention for Codellama-34B and Codellama-7B across varying prompt lengths in the sorting reasoning task. Radix Attention is the best baseline in end-to-end latency.
@@
Wall Clock Time Speedup 7B 1.09$\times$ 1.37$\times$ 1.53$\times$
@@
Table 19: [Different GPUs] Speedup of DeFT in average attention latency (second) with NVIDIA RTX 4090 (24GB) for LLama3-8B model(GQA). Radix Attention is the best baseline in end-to-end latency.
@@
Attention Speedup over the best wall-clock 1.44$\times$ 1.10$\times$ 2.40$\times$
@@
Table 20: [Different Model Architectures(GQA)] Speedup of DeFT in average attention latency (second) with NVIDIA A100(80GB) for Codellama-34B model(GQA). Radix Attention is the best baseline in end-to-end latency.
@@
Attention Speedup over the best wall-clock 1.75$\times$ 1.13 $\times$ 3.37$\times$
Table 21: [Different Model Architectures(MHA)] Speedup of DeFT in average attention latency (second) with NVIDIA A100(80GB) for Codellama-7B model(MHA). Radix Attention is the best baseline in end-to-end latency.
@@
Attention Speedup over the best wall-clock 1.50$\times$ 1.23$\times$ 2.65$\times$

### v4

@@ A.3 Discussion of Concurrent Works
There are some concurrent works (Athiwaratkun et al., 2024; Ye et al., 2024a; Juravsky et al., 2024; Zhu et al., 2024) in attention algorithm design for single-context large-batch sampling, where the goal is to generate multiple sequences from a single context(e.g. system prompt or few-shot examples), which is a special case of tree-based decoding with a depth of 1. The design of their algorithms are based on this feature, which means they can not suit well in attention calculation of a tree with more than two levels of prefixes with efficiency.
@@
Table 9: Comparison among DeFT and concurrent works in single-context large-batch sampling scenarios, including Chunk-Attention (Ye et al., 2024a), Hygragen (Juravsky et al., 2024) and Bifurcated-Attention (Athiwaratkun et al., 2024). RelayAttention (Zhu et al., 2024) and Cascade-inference (Ye et al., 2024b) are similar to Hygragen. More $\star$ means more balanced workloads after tree split, which also shows how insensitive the acceleration is to the tree topology.
@@ A.5 Analysis: IO Complexity of DeFT
Figure 11: (Supplementary to Figure 3) QKV partitioning of Tree Attention (Cai et al., 2024; Miao et al., 2023) and memory access. Tree Attention-Medusa (Cai et al., 2024) partitions QKV by General Matrix Multiply(GEMM) in PyTorch. Tree Attention-SpecInfer (Miao et al., 2023) adopts Q-Guided Grouping. Q-$BCM$ is the Q-Guided bit causal mask for SpecInfer, where each bit represents the causal relationship between a query and a segment of tokens in the KV cache. For example,Q-$BCM$ for $Q_{a}$ is "110" which means the first two segments of KV cache $KV_{0}$ and $KV_{1}$ is valid for $Q_{a}$’s attention. The $Q_{i}$ and $K_{j}$ in the figure are the same as the ones in Figure 3.
@@ A.6 Discussion of Workloads Generation
  • We can utilize decoding trees from real tasks as a benchmark within a unified system, enabling fair comparison of different attention algorithms in terms of decoding latency. This comparison is possible despite the algorithms being based on distinct systems, such as variations in memory management implementations for their kernels.
@@ A.7 Additional Results
The details of attention latency and IO comparison among DeFT and baselines are in Table 5 and Table 17, respectively. Note that the attention latency does not include the memory management overheads, which is the bottleneck for Tree Attention-Medusa. For example, it takes more than $80\%$ of end to end latency in speculative decoding with 32 queries.
@@
Table 16: Average attention latency (in seconds) for tree-based decoding and its impact on decoding latency. Here, $b$ represents the tree width, and $t$ denotes the token tree size (i.e., the number of tree-structured queries). Attention Speedup over the best attention refers to the speedup of DeFT-Flatten compared to the best baseline (typically Tree Attention-Medusa) in attention calculation. Radix Attention is the best baseline for decoding latency. Note that KV cache management is not included in the attention latency. $\star$ denotes out-of-memory (OOM) errors for the A100 80GB GPU. For more details on decoding latency, see Table 5.
@@
Tree Attention-Medusa 3.93 7.51 9.57 38.64 29.10 2.62 3.96 22.40 26.31 41.10 68.28
Paged Radix Attention 5.99 7.30 9.96 39.37 24.69 3.11 5.13 25.73 40.47 76.10 145.43
DeFT-Flatten . 3.47 4.07 5.87 28.41 21.45 2.57 3.83 13.15 16.79 24.46 40.56
Attention Speedup over the best attention. 1.13$\times$ 1.63$\times$ 1.70$\times$ 1.36$\times$ 1.15$\times$ 1.02$\times$ 1.03$\times$ 1.70$\times$ 1.57×1.57$\times$1.57 × 1.68×1.68$\times$1.68 × 1.68×1.68$\times$1.68 ×
Attention Speedup over Radix Attention 1.73$\times$ 1.63$\times$ 1.70$\times$ 1.39$\times$ 1.15$\times$ 1.21$\times$ 1.34$\times$ 1.96$\times$ 2.41×2.41$\times$2.41 × 3.11×3.$1$ 3.59×3.59$\times$3.59 ×
Decoding Speedup over Radix Attention 1.24$\times$ 1.28$\times$ 1.33$\times$ 1.10$\times$ 1.03$\times$ 1.03$\times$ 1.05$\times$ 1.29$\times$ 1.50×1.50$\times$1.50 × 1.91×1.91$\times$1.91 × 2.23×2.23$\times$2.23 ×
Table 17: Average end-to-end IO (TB) during decoding. Data format is Left/Right: (Left) KV Cache IO; (Right) partial results IO, including ${\bm{Q}}{\bm{K}}^{T}$,${\bm{Q}}{\bm{K}}^{\top}/{s_{c}}$, Mask $M$, ${\bm{M}}+{\bm{Q}}{\bm{K}}^{\top}/{s_{c}}$ and Softmax. $b$ means tree width. $t$ denotes the token tree size (i.e., the number of tree-structured queries)$\star$ means out of memory for A100 80GB.
@@
We observe that the effectiveness of attention speedup varies with different decoding tree topologies. Considering the simplest tree structure, a prompt with several suffixes—given a prompt that is not very short, one of the most important factors for speedup is the extent to which we can reuse its KV cache IO. This can be measured by the width of the tree. More specifically, it is determined by the number of queries per iteration. Therefore, we fix the prompt length at 4000 and vary the width of the decoding tree in few-shot prompting (which also indicates how many requests share the same prompt). Then, as shown in Figure 14, we evaluate DeFT-Flatten with the best baseline in attention calculation– Tree Attention-Medusa (Cai et al., 2024) (Medusa-Attn in the figure), as well as the best baseline in decoding latency– Radix Attention (Zheng et al., 2023), for the per-iteration latency over time. We have the following observations:
@@
    Since the tree topology consists of a fixed prefix with several suffixes, a larger tree width allows the prompt prefix’s KV cache to be reused more frequently during IO. This leads to a more significant decoding speedup—1.24×1.24$\times$1.24 × with a width of $w=20$, and 1.33×1.$3$ with a width of $w=50$—compared to Radix Attention.
@@
Figure 14: Per iteration latency for few-shot prompting tasks with different tree width. e2e means decoding latency(optimal end-to-end latency), while Attn means only the attention overhead.
@@
Ablation: The influence of prompt length.
We set the prompt length from 1k to 20k tokens, and the generation length is set to 1k tokens, to simulate a long prompt short generation case. Our workloads are speculative decoding with token tree size of 32 and 64. The ablation is based on Llama3.1-8B model and NVIDIA A100 80GB. We have three metrics: time per output token (Figure 16), decoding latency (Figure 17) and attention latency (Figure 18).
Figure 16: Time per output token(TPOT) of DeFT with different prompt lengths in speculative decoding.
Figure 17: Decoding latency of DeFT with different prompt lengths in speculative decoding.
Figure 18: Attention latency of DeFT with different prompt lengths in speculative decoding.
@@
In Table 6 and Table 8 of Section 4.4, we show the influence of prompt length and model size, individually. Here, we present the ablation study of both model size and prompt length for sorting task, as shown in Table 18. With a longer prompt, DeFT shows more pronounced speedup in the same model, since the attention overhead is proportional to the token count in the decoding tree, while the FFN overhead remains nearly constant for the same model. For a fixed prompt length, with the larger model—Codellama-34B, DeFT-Flatten achieves slightly reduced but still significant (up to 1.28x) decoding speedup. The performance reduction is attributed to a lower A/F-LR, as the FFN overhead is greater in larger models.
Table 18: [Ablation Study of Model Size and Prompt Length] Comparison of decoding speedup and Attention/FFN latency ratio (A/F-LR) between DeFT and Radix Attention for Codellama-34B and Codellama-7B across varying prompt lengths in the sorting reasoning task. Radix Attention is the best baseline in decoding latency.
@@
Decoding latency Speedup 7B 1.09$\times$ 1.37$\times$ 1.53$\times$
@@
Table 19: [Different GPUs] Speedup of DeFT in average attention latency (second) with NVIDIA RTX 4090 (24GB) for LLama3-8B model(GQA). Radix Attention is the best baseline in decoding latency.
@@
Attention Speedup over the best decoding 1.44$\times$ 1.10$\times$ 2.40$\times$
@@
Table 20: [Different Model Architectures(GQA)] Speedup of DeFT in average attention latency (second) with NVIDIA A100(80GB) for Codellama-34B model(GQA). Radix Attention is the best baseline in decoding latency.
@@
Attention Speedup over the best decoding 1.75$\times$ 1.13 $\times$ 3.37$\times$
Table 21: [Different Model Architectures(MHA)] Speedup of DeFT in average attention latency (second) with NVIDIA A100(80GB) for Codellama-7B model(MHA). Radix Attention is the best baseline in decoding latency.
@@
Attention Speedup over the best decoding 1.50$\times$ 1.23$\times$ 2.65$\times$